### v1

# Controllable Patching for Compute-Adaptive Surrogate Modeling of Partial Differential Equations

Abstract
Patch-based transformer surrogates have become increasingly effective for modeling spatiotemporal dynamics, but the fixed patch size is a major limitation for budget-conscience deployment in production. We introduce two lightweight, architecture-agnostic modules—the Convolutional Kernel Modulator (CKM) and Convolutional Stride Modulator (CSM)—that enable dynamic patch size control at inference in patch-based models, without retraining or accuracy loss. Combined with a cyclic patch-size rollout, our method mitigates patch artifacts and improves long-term stability for video-like prediction tasks. Applied to a range of challenging 2D and 3D PDE benchmarks, our approach improves rollout fidelity and runtime efficiency. To our knowledge, this is the first framework to enable inference-time patch-size tunability in patch-based PDE surrogates. Its plug-and-play design makes it broadly applicable across architectures—establishing a general foundation for compute-adaptive modeling in PDE surrogate tasks  [^1] Full codebase will be made publicly available upon publication in this repository..
1 Introduction
Numerical methods have long been the gold standard for simulating spatio-temporal dynamics in systems governed by partial differential equations (PDEs). Numerical solvers offer convergence guarantees, exploit problem sparsity, and provide fine-grained control of the accuracy-compute trade-off through configurable resolution [1, 2]. This last aspect is of vital importance as it is well-known that certain aspects of physics can only be resolved with sufficient resolution so that the choice of resolution is driven by both physical necessity and the availability of compute[3, 4].
Figure 1: Overview diagram of CKM and CSM approaches for adaptive patching in ViT-based PDE surrogate models. Here, $k$ and $s$ denote the convolutional kernel size and stride, respectively, which control the downsampling rate. Additionally, this figure highlights our rollout strategy at inference, where the dynamic kernel or stride alternates cyclically between 4, 8, and 16 at each timestep.
In recent years, deep learning and particularly transformers have emerged as a powerful framework for training surrogate models [5, 6, 7, 8]. Often derived from computer vision methods such as the Vision Transformer (ViT [9]), these models segment discretized fields into distinct, non-overlapping patches, converting them into tokens for the transformer. This results in a downsampling of the input, an essential step as self-attention cost scales quadratically with token count. While initially expensive to train, this cost can be amortized by repeated usage in problems such as forecasting [10], PDE-constrained optimization [11], and parameter inference [12, 13].
However, the inflexible computational cost of the ViT-based framework can become an undue burden as models are deployed in production. At a fixed input resolution, a ViT-based model incurs a fixed cost, regardless of problem complexity or the budget of the user. This problem is not unique to physical simulation - notably, NLP researchers have explored variable-cost solutions, such as switching between models of different expense [14], adjusting hidden dimension size [15], or other test-time scaling methods [16].
While the problem is not unique to physical settings, the physics domain does offer more natural paths toward adaptive compute. It is well understood in physical modeling that there are often hard resolution thresholds required for resolving fine-scale features like shocks or wave-fronts [17]. In many problems, these small scales can influence larger scales overtime and thus resolving them can lead to significant accuracy gains. For ViT-style models, the internal resolution is entirely determined by the patch size used for tokenization. This can influence both model accuracy and computational cost [9, 18]. Smaller patches act to reduce the level of downsampling, allowing the transformer to resolve finer details with greater accuracy (Fig. 3) but at significant cost. Thus, we seek a compute-adaptive approach for balancing the accuracy and budget demands of users by controlling the model’s internal patch resolution.
To address this issue, this paper introduces two different strategies for adjusting downsampling rates in patch-based tokenization blocks enabling controllable resolution transformers and giving users more precise control over their desired location on compute-accuracy frontier. Convolutional Kernel Modulation (CKM) utilizes the patch-resizing of FlexiViT [19] in both the encoder and decoder while Convolutional Stride Modulation (CSM) utilizes a static convolutional kernel but adjusts the compression rate directly by modulating the stride. These two strategies are visualized in Fig. 1.
Crucially, this flexibility opens up new modeling capabilities that are simply unavailable to static-patch architectures. For example, we show that during autoregressive rollout — where errors can accumulate over time — alternating patch sizes across timesteps mitigates patch-induced artifacts such as harmonic spectral spikes (see Fig. 2). This artifact reduction highlights the emergent benefits of enabling patch/stride diversity at inference, leading to broader implications for stability and robustness in long-horizon surrogate modeling.
Figure 2: Left: Residual power spectrum of the density field at the 20th rollout step for CKM, CSM, and fixed patch size 16 (p16) on the turbulent_radiative_layer_2D dataset [20]. Harmonic spikes indicate patch-induced artifacts, significantly reduced in CKM and CSM. Right: Prediction at rollout step 20, where CKM and CSM mitigate visible patching effects in fixed patch 16 (P16) model.
Contributions.
We propose a general framework for enabling compute-adaptive inference in transformer-based PDE surrogates. Our contributions span model design, architectural generality, and empirical insights:
  1. 1.
    Controllable Tokenization: We introduce Convolutional Kernel Modulation (CKM) and Convolutional Stride Modulation (CSM) as two approaches for controllable tokenization. Used with randomized compression rates during training, they allow test-time compute-adaptive accuracy.
  2. 2.
    Architecture-agnostic deployment: These flexible patching and striding strategies operate at the tokenization level of the encoder and decoder, and are compatible with a broad class of transformer architectures. We show this generality using two different types of standard base transformer processors – vanilla ViT [9] with full attention and axial ViT [21].
  3. 3.
    Competitive to fixed patch size baselines: Across a diverse suite of 2D and 3D PDE datasets from the Well collection [20], CKM and CSM consistently match or outperform fixed-patch size surrogates at multiple token counts—despite being trained once. This holds for both next-step prediction and long autoregressive rollouts. Our models are also shown to be competitive with a range of state-of-the-art non-patch-based baselines, underscoring their broader importance.
  4. 4.
    Patch Artifact Reduction: Leveraging the CKM and CSM models mentioned above, we introduce a novel rollout strategy that mitigates patch-induced artifacts (spikes at harmonics of the inverse patch size, see Fig. 2). This significantly enhances the quality of long-term predictions, especially in capturing finer-scale physical features. This artifact reduction emerges naturally from the ability to modulate tokenization over time—demonstrating an additional benefit of flexible inference.
2 Background and related work
Notations. We focus on autoregressive surrogate models for time-evolving PDE systems, where the goal is to predict the next state of a discretized physical field from a short temporal context. This setup mirrors video prediction on a uniform grid. We refer to our proposed CKM/CSM models as compute-elastic, meaning they support tunable patch or stride resolution at inference time. Unless otherwise stated, scalability refers to this token-level flexibility, not mesh-level resolution changes.
Architecture-agnostic tokenization for PDE surrogates. CKM and CSM are architecture-agnostic tokenization strategies applied at the encoder and decoder stages. They control the patch/stride configuration for spatial downsampling while remaining independent of the transformer architecture. We demonstrate compatibility with standard vanilla ViTs (full attention) [9] and axial ViTs [22]. This modularity makes CKM and CSM flexible plug-ins for a wide range of transformer-based PDE surrogates, enabling adaptive deployment without redesign. To our knowledge, this is the first work to propose such flexible, architecture-agnostic tokenization for PDE surrogates.
Surrogate modeling of PDEs with deep learning. PINNs [23, 24, 25, 26, 27, 28, 29, 30, 31] embed physics in loss functions, while CNN- and GNN-based surrogates operate on structured or unstructured meshes, supporting spatiotemporal tasks such as climate and fluid modeling [32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46]. Neural operators [47, 48, 49, 50, 51, 52, 53, 5, 54, 55, 56, 57, 58] aim for resolution-invariant learning, with FNOs [59] using Fourier transforms for global context.
ViTs and compute adaptive models for PDE surrogate modeling. Initially developed for image recognition tasks, ViTs [9] have seen increasing adoption in the PDE surrogate modeling literature [60, 6, 61]. Despite their strengths, traditional ViTs rely on fixed patch sizes, which strongly limits their flexibility and adaptability when modeling PDE–based systems as discussed in the introduction. Adaptive patching enables dynamic adjustment of patch sizes based on task complexity, as seen in FlexiViT [19] for vision tasks, which adapts ViTs to varying patch sizes and compute budgets. Although explored in computer vision, adaptive patching for PDE surrogate modeling has, to our knowledge, only been explored in [62], the difference with our work being that their patchification is data-driven, while ours can be manually controlled based on compute requirements.
3 Flexible encoder & decoder frameworks
Overview. CKM and CSM operate as encoder-decoder modules within a ViT based architecture for autoregressive PDE prediction. The transformer core applies a sequence of temporal attention, spatial attention, and MLP blocks (App. B.2). Crucially, our framework is architecture-agnostic—we test it with axial attention and vanilla ViTs (and also a recent model– Continuous ViT [63] in App. D).
Utility of compute-elastic models. As illustrated in Fig.3, increasing the number of tokens—equivalent to using smaller patch sizes in models with fixed patches—enhances the accuracy of autoregressive prediction tasks, but at a higher computational cost. The cost for models trained with fixed patch sizes cannot be controlled at inference time. This limits the utility of these models in budget-aware production settings and motivate the exploration of compute-elastic models for PDE modeling with ViT-inspired architectures.
Research in the broader machine learning community has extensively demonstrated the value of scale [64, 65]. As transformer-based surrogate models similarly scale in parameter count, ensuring inference-time scalability is crucial, particularly to ensure these models are able to be used in compute-constrained environments. To avoid any confusion, by inference scalability, we mean patching/striding scalability or tunability, and not scalability in mesh resolution. Training and maintaining multiple fixed-patch models is a burden especially as very large models, such as large pretrained or foundation models are being built [7, 8, 66]. In such a scenario, if our compute-elastic strategies can enable training of a single model without retraining for multiple patch resolutions without any loss in accuracy, these strategies can form the building blocks for larger scale foundation models for PDEs.
Patching.
Given an input image with height $H$ and width $W$, ViTs will divide the image into non-overlapping patches of size $k\times k$, producing $N=\lfloor H/k\rfloor\cdot\lfloor W/k\rfloor$ tokens. For global attention, transformer cost scales as $\mathcal{O}(N^{2})$, and for axial attention, as $\mathcal{O}(N\sqrt{N})$ [22]. In convolutional tokenization, patches are extracted via kernels of size $k$ and stride $s$, resulting in a token count
$N\coloneqq N_{h}\cdot N_{w},\quad\text{for}\quad N_{h}\coloneqq\left\lfloor\frac{H-k}{s}\right\rfloor+1,\quad N_{w}\coloneqq\left\lfloor\frac{W-k}{s}\right\rfloor+1,\quad$ (1)
While most ViT based architectures use $k=s$, our CKM and CSM modules decouple these parameters to flexibly control token count at inference.
Convolutional Kernel Modulator (CKM).
CKM introduces dynamic patch size selection within convolutional encoder and decoder blocks, enabling a model to tokenize input fields using patch sizes $k$in {4, 8, 16} at each forward pass. These powers-of-two sizes reflect standard discretizations in PDE data, and contrast with most ViT-based surrogates that fix patch size at 16 throughout training and inference [7, 66].
Consider input $x\in\mathcal{X}^{T}$ for some input space $\mathcal{X}^{T}\subseteq\mathbb{R}^{M\times H\times W\times T\times C}$, with $M,T$ and $C$ being the batch size, time context and input channels respectively. CKM resizes a base convolutional kernel $w^{\text{base}}\in\mathbb{R}^{k_{\text{base}}\times k_{\text{base}}\times C\times C^{\prime}}$ using PI–resize [19] (App. A):
$w=B^{\mathsf{T}\dagger}w^{\text{base}},\quad x_{\text{enc}}=\operatorname{Conv}(x,w,\texttt{stride}=k),$ (2)
where $B\in\mathbb{R}^{k\times k_{\text{base}}}$ is a bicubic interpolation matrix and $B^{\mathsf{T}\dagger}$ its pseudoinverse. Tokens are processed by the transformer processor (this transformer can have any architecture), producing latents $x_{\text{lat}}\in\mathcal{X}_{\text{lat}}$ for some latent space $\mathcal{X}_{\text{lat}}$. The decoder reuses the same $k$ and $w$ to reconstruct:
$\hat{x}=\operatorname{ConvTranspose}(x_{\text{lat}},w,\texttt{stride}=k),$ (3)
yielding a next-step prediction $\hat{x}\in\mathcal{X}^{1}\subseteq\mathbb{R}^{B\times H\times W\times 1\times C}$. Pseudocode is in App. A.5.
Comparison with FlexiViT. While FlexiViT introduced flexible patching in image classification via a single-stage encoder [19], it lacked a decoder and was not designed for spatiotemporal forecasting. CKM generalizes this idea, enabling PI–resize modulation across any convolutional stage—encoder or decoder, single- or multi-stage—as long as convolutional weights are available. FlexiViT is thus a special case within the broader CKM framework.
Modular Applicability. In our vanilla ViT and Axial ViT models, we adopt two-stage convolutional encoders and decoders following the hMLP architecture [67], which has recently been adopted for large-scale PDE surrogates [66, 7], and apply CKM independently at each stage.
Convolutional Stride Modulator (CSM).
CSM enables flexible downsampling by sampling the stride $s\in\{4,8,16\}$ at each forward pass. Given input $x\in\mathcal{X}^{T}$, we pad the input with learned tokens (informed by boundary conditions) to avoid edge artifacts. Using a fixed kernel $k^{\text{base}}$, the encoder produces:
$x_{\text{enc}}=\operatorname{Conv}(x,k^{\text{base}},\texttt{stride}=s),$ (4)
which is passed through a transformer to obtain $x_{\text{lat}}$. The decoder applies a transposed convolution:
$\hat{x}=\operatorname{ConvTranspose}(x_{\text{lat}},k^{\text{base}},\texttt{stride}=s),$ (5)
to generate the autoregressive prediction $\hat{x}\in\mathcal{X}^{1}$. Like CKM, CSM is compatible with both single-stage and multi-stage patching schemes. Together, CKM and CSM offer general strategies for flexifying convolutional tokenization in ViTs. Pseudocode is in App. A.6.
Novel rollout strategy.
In traditional autoregressive forecasting, patch size is fixed across time. Our compute-elastic modules unlock a novel strategy: alternating patching during rollout. At each step, the model cycles through $k\in\{4,8,16\}$ for CKM (or $s\in\{4,8,16\}$ for CSM). This test-time scheduling is entirely decoupled from training and enables new temporal patterns in tokenization.
Surprisingly, we find this alternating rollout suppresses patch artifacts—periodic errors that appear as harmonics in the spectral residuals (see Fig. 2). These artifacts are ubiquitous in fixed-patch ViT based architectures, regardless of attention type (vanilla, axial, Swin), and cannot be removed through training alone (App. E). Alternating rollouts yield cleaner, more stable long-horizon predictions. Importantly, this discovery was only possible due to the test-time flexibility provided by our framework—a capability entirely absent from prior PDE surrogates.
4 Experimental Results
To evaluate the performance of our compute-elastic models, we conduct experiments on 2D and 3D datasets from The Well [20]. App. B.1 for specifics. The experiments are briefly described below:
  1. 1.
    We compare CKM and CSM to SOTA non-patch based PDE surrogates including TFNO [68], FFNO [69], SineNet [70], and Transolver [71]. These comparisons demonstrate that flexible patch-based models are not only viable but often outperform non-tokenized architectures in next-step and rollout prediction tasks.
  2. 2.
    We test whether a single compute-elastic model is preferable to training multiple fixed-patch models. We train three fixed-patch models (patch sizes 4, 8, 16) for 200 epochs each (App. B.3), then train CKM and CSM models using the combined compute budget.
    This setup reflects a practical question: given a fixed training budget, is it better to train one flexible model or several static ones? All evaluations are performed on models trained under this constraint. After training, we compare the performance of CKM and CSM models against individual models trained at static patch sizes. If the compute elastic models can match or improve over the fixed patch baselines, this enables the novel creation of a broad class of flexible PDE surrogate eliminating the need to train and maintain separate fixed-patch models.
  3. 3.
    We perform the experiments described above on two transformer-based processors, namely the axial ViT [22] and the vanilla ViT [9] to demonstrate the architecture agnostic–ness of the CKM/CSM strategies. [9].
  4. 4.
    We assess how these flexible modules improve stability, prediction accuracy and mitigate artifacts in extended rollouts (Sec.4.3).
We additionally provide a set of ablation studies in App G. The input spatio-temporal sequence consists of six time frames, with the loss function optimized for next-step prediction.
Figure 3: Next-step prediction test VRMSEs vs number of tokens at inference of a 100M parameter model trained on four 2D datasets (left to right):shear_flow, turbulent_radiative_layer_2D, active_matter and  rayleigh_benard. Lower VRMSE is better. Token count is a proxy for required compute. This plot highlights the compute-accuracy tradeoff expanded more in sec. 4.2
4.1 Comparison to non-patch based SOTA models
Although our focus is not on non-patch architectures, Table 1 provides a reference comparison to state-of-the-art CNN/Neural-Operator surrogates. It demonstrates that flexible models are not only viable, but highly competitive with SOTA PDE surrogate architectures that do not use patch-based tokenization (e.g., TFNO [68], FFNO [69], Transolver [71], SineNet [70]). In fact, our models often significantly outperform these baselines. This result reinforces the potential of adaptive patch-based transformers to serve as competitive surrogates for PDE modeling, without sacrificing accuracy or flexibility. We did this architectural comparison for two complex physical systems, shear_flow  and turbulent_radiative_layer_2D  from the Well [20]. In the results of the sections to follow, we will focus mainly on comparisons between our flexible models vs. static patch baselines.
Model TFNO FFNO SineNet Transolver CSM + ViT (Ours) CKM + ViT (Ours)
Number of parameters 20M 7M 35M 11M 7M 100M 7M 100M
shear_flow 0.89 0.11 0.17 $>1$ 0.0762 0.0538 0.096 0.0531
turbulent_radiative_layer_2D 0.81 0.485 0.65 $>1$ 0.458 0.299 0.477 0.321
Table 1: 10-step rollout VRMSE comparison across patch-free models and our patch-based surrogate variants. We report scores for both 7M and 100M parameter versions of CSM and CKM.
4.2 Prediction accuracy and inference time patch/stride tunability
Dataset # Tokens Axial ViT (50M) Vanilla ViT (100M) The Well Baselines
CKM CSM f.p.s. CKM CSM f.p.s.
shear_flow 2048 0.0070 0.0070 0.0085 0.0055 0.0055 0.0067 CNext-U-Net
512 0.0098 0.0104 0.0124 0.0086 0.0088 0.0096 0.1049
128 0.0143 0.0134 0.0173 0.0120 0.0110 0.0140
turbulent_radiative_layer_2D 3072 0.153 0.178 0.170 0.133 0.144 0.143 CNext-U-Net
768 0.165 0.178 0.186 0.153 0.178 0.170 0.2269
192 0.211 0.211 0.237 0.200 0.202 0.223
active_matter 4096 0.0240 0.0230 0.0275 0.0192 0.0172 0.0210 CNext-U-Net
1024 0.025 0.028 0.033 0.0220 0.0210 0.0240 0.0330
256 0.0369 0.0367 0.0430 0.034 0.029 0.031
rayleigh_benard 4096 0.0277 0.0290 0.0361 0.0250 0.0250 0.0310 CNext-U-Net
1024 0.0370 0.0398 0.0442 0.0330 0.0350 0.0360 0.2240
256 0.0538 0.0522 0.0581 0.0470 0.0460 0.0470
supernova_explosion 4096 0.288 0.305 0.288 0.266 0.287 0.272 U-Net
512 0.370 0.380 0.392 0.360 0.380 0.387 0.3063
64 0.419 0.421 0.434 0.413 0.417 0.426
turbulence_gravity_cooling 4096 0.124 0.138 0.150 0.103 0.121 0.118 CNext-U-Net
512 0.180 0.197 0.194 0.164 0.182 0.179 0.2096
64 0.216 0.215 0.214 0.202 0.200 0.198
Table 2: Test VRMSE across multiple 2D and 3D PDE datasets for next-step prediction using CKM, CSM, and fixed patch size (f.p.s.) models. CKM/CSM models are trained once and evaluated across multiple token counts (i.e., patch/stride configurations), whereas each fixed p.s. model is trained separately. Results are shown for two transformer backbones: Axial ViT (50M) [22] and Vanilla ViT (100M) [9]. We also report the best outcomes from the Well benchmark suite [20].
We choose the VRMSE metric (App. B.4) as a measure of accuracy. In Table 2 and Fig. 3, we report VRMSE across a range of 2D and 3D PDE datasets from the Well [20]. Each model is evaluated at multiple token counts, which correspond to varying patch sizes in CKM or stride sizes in CSM —highlighting the impact of inference-time compression on predictive accuracy. Importantly, the fixed patch size (p.s.) models are separate networks trained separately at each patch size, while CKM/CSM are trained only once using random patch/stride schedules and evaluated without retraining. We show results for two different transformer architectures, Axial ViT [22] and vanilla full attention ViT [9]. Importantly, the goal is not to compare the two base architectures against each other, but to show that for both architectures, higher token count implies greater accuracy, and that CKM/CSM can enable a token count tunability at inference against the corresponding fixed patch baselines without losing any accuracy.
Across all datasets and token configurations, CKM and CSM consistently match or outperform their fixed-patch counterparts—despite being single, compute-elastic models. This includes both axial and vanilla attention processors, demonstrating the architecture-agnostic nature of our encoder/decoder modifications. Across all datasets, we observe a consistent trend: increasing the number of tokens at inference—by using smaller patch sizes (CKM) or strides (CSM)—leads to improved VRMSE, but it is more expensive computationally. This behavior is shown in both Table 2 and Fig.3. For example, in the active_matter  dataset, with a spatial resolution of $256^{2}$, reducing the patch/stride size from 16 to 4 increases token count from 256 to 4096, triples inference time (0.21s/step → 0.63s/step), and increases compute from 5 to 170 GFLOPs—but also reduces error by over 30%. Similarly, for our 3D datasets with a spatial resolution $64^{3}$, decreasing the patch size from 4 to 16, or increasing the tokens from 64 to 4096 leads to an increase in the inference speed per step by approximately a factor of 8 (0.11s/step→ 0.8s/step); but also improves VRMSE by a factor of $\sim$2.
These results reinforce that token count serves as a natural proxy for compute-accuracy tradeoff in patch-based PDE surrogates. Importantly, CKM and CSM allow a single model to be flexibly deployed at different compute budgets, without retraining, unlike fixed-patch baselines. We additionally evaluate performance across spatial frequencies using the binned spectral normalized mean squared error (BSNMSE). Across datasets, CKM and CSM consistently match or outperform fixed patch models in all frequency bands, highlighting their ability to preserve both coarse and fine-scale features. Further definitions, and visualizations are provided in App.  C.
Finally, as also stated before, while our main goal is not to benchmark against non-patch-based surrogates, Table 2 still includes the best results from FNO [72], TFNO [68], U-Net, and CNext-U-Net [73] from the Well benchmark suite [20]. We re-trained these models with tuned learning rates and report their best accuracy. Despite this, our flexible models consistently outperform these baselines across most tasks and inference budgets, demonstrating that flexible patch-based surrogates are not only viable but can exceed the performance of traditional operator-based or convolutional surrogates, and outperform a range of SOTA architectures on complex datasets as discussed in sec. 4.1.
To further test the modularity of our flexible modules, we integrate it into the encoder of the Continuous Vision Transformer (CViT) [63], a recent hybrid architecture which features a patch-based transformer encoder coupled with a novel, grid- based continuous decoder that directly queries spatiotemporal locations. Since the decoder in CViT does not use patch-based tokenization, we apply CKM only to the encoder. Since CViT is designed for 2D problems, we restrict our evaluation to 2D datasets, ensuring a fair and focused test of patch flexibility without altering any other model components. Results across four 2D PDE datasets show that a flexified CViT consistently outperforms its fixed-patch CViT counterpart for both next–step and rollout results. This experiment underscores that CKM can be selectively applied even within hybrid ViT models, further reinforcing the architecture agnostic-ness and modularity of these strategies. Full results are presented in App.D.
This ability to dynamically scale computational resources at inference is, to our knowledge, new in the PDE emulator literature. While large language models leverage inference-time optimizations such as entropy-based techniques, PDE surrogate models typically lack such flexibility. Our compute-elastic models address this gap by enabling adaptive patches and strides, allowing dynamic resource allocation to balance accuracy and efficiency at inference.
Figure 4: Step 44 rollout for the $D_{xy}$ field of the active_matter dataset.
4.3 Improved rollouts with compute elastic models
A key advantage of compute-elastic models is their ability to enable novel inference-time rollout strategies that are inaccessible to fixed-architecture surrogates. We introduce one such strategy: an alternating-patch rollout, where CKM and CSM models dynamically cycle through patch or stride sizes of 4, 8, and 16 at each prediction step. To our knowledge, this is the first demonstration in the PDE surrogate modeling literature where inference-time patching is not only treated as a tunable knob, but actively used to improve long-horizon prediction fidelity.
This strategy significantly reduces a persistent issue in patch-based architectures: patch artifacts—grid-like distortions that commonly arise in autoregressive rollouts, especially for fixed patching. These artifacts, prevalent across ViT variants (vanilla, axial, Swin), stem not from training instability or data quality, but from the limitations of static tokenization (App. E for more extended evidence).
Figure 4 demonstrates this phenomenon visually in the active_matter  dataset. We see significant checkerboard artifacts at rollout step 44 for the static patch model, which are largely stabilized by our alternating rollout strategy. Quantitatively, Table 3 shows the 10-step rollout VRMSE for all datasets. Across both axial and vanilla transformer backbones, compute-elastic models consistently outperform fixed patch models—often by large margins—demonstrating greater long-horizon stability. In fact, comparing these 10-step rollout values again with the non-patch SOTA models in Table 1 for relevant datasets, we find that from a purely architectural comparison perspective, our models often achieve SOTA outcomes on these complex tasks. Fig. 5 additionally shows a full trajectory evolution of the shear_flow dataset comparing fixed patch vs CKM/CSM. Clearly, the fixed patch 16 model develops prominent artifacts already by step 40, while CKM/CSM (CKM especially in this case) maintain a much more stable long trajectory. For additional rollout visualizations and metrics, refer to the App. F. We emphasize the powerful insight here that these improvements emerge without any additional training, enabled purely through dynamic inference-time patch modulation.
Figure 5: 100 step full trajectory rollout of shear_flow . Top to bottom: ground truth, fixed patch 16 model, CKM+vanilla ViT and CSM+vanilla ViT.
Dataset Axial ViT (50M) Vanilla ViT (100M)
CKM CSM f.p.s. $\Delta$ CKM CSM f.p.s. $\Delta$
shear_flow 0.082 0.089 0.143 +42.7% 0.057 0.053 0.107 +50.5%
turbulent_radiative_layer_2D 0.570 0.475 0.571 +16.8% 0.409 0.373 0.446 +16.4%
active_matter 0.390 0.384 0.640 +40.0% 0.351 0.359 0.370 +5.1%
rayleigh_benard 0.2159 0.166 0.217 +23.5% 0.215 0.140 0.2273 +38.4%
supernova_explosion 1.270 1.310 1.905 +33.3% 1.14 1.20 1.75 +34.9%
turbulence_gravity_cooling 0.680 0.667 0.860 +22.4% 0.527 0.559 0.77 +31.6%
Table 3: 10-step rollout VRMSE comparison across datasets for CKM/CSM vs fixed patch size (f.p.s) models. $\Delta$ indicates the best percentage improvement (i.e., lowest VRMSE vs. fixed patch baseline) for each transformer variant.
5 Ablations
We perform ablations to assess the robustness of our approach. Varying model size (7M, 25M, 100M) reveals consistent trends: (1) smaller patches improve accuracy; (2) CKM/CSM, trained once, match or outperform fixed-patch models; and (3) flexibility enables compute-adaptive inference without retraining (App. G.1).
We also ablate the base patch size used in PI-resize projection. Among ${4,8,16,32}$, the differences are not major (App. G.2). Reducing patch diversity during training (e.g., omitting $P=8$) degrades performance at the excluded size, underscoring the value of patch variety (App. G.3).
Finally, we compare cyclic vs. random patch switching during rollouts. Random schedules significantly hurt performance (App. Table G.4), yet this test-time flexibility underscores a new capability: probing rollout behaviors without retraining—an option simply not available in prior PDE surrogates.
Conclusion
We introduced compute-elastic encoder/decoder modules—CKM and CSM —that allow inference-time control over patch and stride sizes in transformer-based PDE surrogates. Across challenging 2D and 3D benchmarks, these flexible models consistently match or exceed fixed-patch baselines while enabling real-time trade-offs between accuracy and computational cost. This adaptability is critical as surrogate models scale toward general-purpose, deployable architectures.
Beyond accuracy–cost control, token-level modulation enables new rollout strategies. Cyclically varying patch/stride sizes during rollout significantly reduces long-horizon artifacts—a persistent issue in ViT-based surrogates—unlocking capabilities previously unavailable in PDE modeling. Our rollout strategy exemplifies how such control can expose patch-sensitive behaviors, and could pave the way for discovering other dynamics that static architectures obscure.
Future directions. If compute resources were not a limiting factor, our study could be extended along several directions. For example, we could test our CKM/CSM modules for flexifying non-ViT but patch-based architectures, such as AFNO [74], or state-of-the-art neural operator models like DPOT [75], which also benefit from patch-size scaling (App. H). Additionally, our strategy could be applied to other ViT variants like Swin [76] or CSWin [77]. Another promising direction is integrating CKM/CSM into large pretrained or foundation models [7, 8, 66], where inference-time patch control could support downstream finetuning across diverse tasks—each naturally operating at different points along the compute–accuracy frontier.
We anticipate that our flexible patching modules can generalize across ViT and non-ViT backbones alike. Our study provides a foundation for future exploration of flexible, compute-aware modeling strategies in the PDE surrogate literature, and opens the door for building next-generation surrogate models with tunable inference costs, modular components, and improved generalization.
Acknowledgements
The authors would like to thank Géraud Krawezik and the Scientific Computing Core at the Flatiron Institute, a division of the Simons Foundation, for the compute facilities and support. Polymathic AI acknowledges funding from the Simons Foundation and Schmidt Sciences, LLC.
Appendix A Derivation of the PI-Resize Matrix for Convolutional Kernel Modulators
A.1 Motivation for Kernel Resizing
The CKM-based model dynamically adjusts convolutional kernel sizes at each forward pass while maintaining a fixed architecture. This enables variable patch sizes ($4,8,16$), which is crucial for handling different spatial resolutions while ensuring compatibility with the ViT.
However, this flexibility introduces a problem: the CNN encoder, which extracts ViT tokens, is trained with a fixed base convolutional kernel. To ensure that features extracted at different patch sizes remain aligned, we use a resize transformation, which projects the base kernel to a dynamically selected kernel size [19].
A.2 Mathematical Formulation
Let:
  • $W^{\text{base}}\in\mathbb{R}^{k^{\text{base}}\times k^{\text{base}}\times c_{\text{in}}\times c_{\text{out}}}$ be the learned CNN weights for the fixed base kernel size $k^{\text{base}}$.
  • $W\in\mathbb{R}^{k\times k\times c_{\text{in}}\times c_{\text{out}}}$ be the resized kernel for a dynamically selected kernel size $k$.
  • $B$ be an interpolation matrix (e.g., bilinear, bicubic) that resizes input patches from size $k$ to size $k^{\text{base}}$.
  • $x\in\mathbb{R}^{B\times k\times k\times c_{\text{in}}}$ be an input patch, where:
    • $B$ is the batch size.
    • $k\times k$ is the local receptive field defined by the kernel.
    • $c_{\text{in}}$ is the number of input channels.
Intuitively, the goal is to find a new set of patch- embedding weights $W$ such that the tokens of the resized patch match the tokens of the original patch
$W^{\text{base}}\ast x\approx W\ast(Bx),$ (6)
where $\ast$ denotes convolution. The goal is to solve for $W$. Mathematically, this is an optimization problem.
A.3 Least-Squares Optimization for PI-Resize
A.3.1 Expanding the Objective
We aim to expand the expectation:
$\mathbb{E}_{x\sim\mathcal{X}}\left[(x^{T}W^{\text{base}}-x^{T}B^{T}W)^{2}\right].$ (7)
Since inner products can be rewritten as matrix-vector multiplications, we note that:
$\langle x,W^{\text{base}}\rangle=x^{T}W^{\text{base}}.$
Step 1: Expand the Square
Using the identity $(a-b)^{2}=a^{2}-2ab+b^{2}$, we expand:
$(x^{T}W^{\text{base}}-x^{T}B^{T}W)^{2}$ $=(x^{T}W^{\text{base}})^{2}-2x^{T}W^{\text{base}}(x^{T}B^{T}W)+(x^{T}B^{T}W)^{2}.$ (8)
Step 2: Take the Expectation
Now, we take expectation over $x\sim\mathcal{X}$:
$\mathbb{E}_{x\sim\mathcal{X}}\left[(x^{T}W^{\text{base}})^{2}-2x^{T}W^{\text{base}}(x^{T}B^{T}W)+(x^{T}B^{T}W)^{2}\right].$ (9)
Using the definition of the covariance matrix:
$\Sigma=\mathbb{E}_{x\sim\mathcal{X}}[xx^{T}],$ (10)
we apply the linearity of expectation to each term:
$\mathbb{E}_{x}[(x^{T}W^{\text{base}})^{2}]$ $=W^{\text{base},T}\Sigma W^{\text{base}},$ (11)
$\mathbb{E}_{x}[(x^{T}B^{T}W)^{2}]$ $=W^{T}B\Sigma B^{T}W,$ (12)
$\mathbb{E}_{x}[x^{T}W^{\text{base}}(x^{T}B^{T}W)]$ $=W^{\text{base},T}\Sigma B^{T}W.$ (13)
Step 3: Write in Matrix Form
Substituting these into the expectation:
$\mathbb{E}_{x\sim\mathcal{X}}\left[(x^{T}W^{\text{base}}-x^{T}B^{T}W)^{2}\right]$ $=W^{\text{base},T}\Sigma W^{\text{base}}-2W^{\text{base},T}\Sigma B^{T}W+W^{T}B\Sigma B^{T}W.$ (14)
The above expression can be re-written as:
$\|W^{base}-B^{T}W\|^{2}_{\Sigma}=(W^{base,T}-W^{T}B)\Sigma(W^{base}-B^{T}W).$ (15)
Conclusion
This result expresses the squared error between the transformed weight matrix $W$ and the resized weight matrix $B^{T}W^{\text{base}}$, weighted by the covariance matrix $\Sigma$. The quadratic form measures the deviation in terms of feature space alignment, ensuring that the transformation remains consistent with the learned base kernel.
The optimal solution minimizing Eq. 15 is given by
$W=\left(\sqrt{\Sigma}B^{T}\right)^{\dagger}\sqrt{\Sigma}W^{base}$ (16)
where $\left(\sqrt{\Sigma}B^{T}\right)^{\dagger}\sqrt{\Sigma}$ is the Moore-Penrose pseudoinverse matrix.
A.4 Final Transformation and Implementation
Thus, at each forward pass, we apply the following steps:
  1. 1.
    Randomly select a kernel size $k\in\{4,8,16\}$.
  2. 2.
    Compute the interpolation matrix $B$ that resizes the input patch from $k$ to $k^{\text{base}}$.
  3. 3.
    Transform the kernel weights using:
    $W=(B^{T})^{\dagger}W^{\text{base}}.$ (17)
  4. 4.
    Apply the resized kernel $W$ for convolution.
This ensures that dynamically changing kernel sizes maintain compatibility with the fixed architecture while enabling flexible patch sizes.
A.5 CKM pseudocode
Algorithm 1 Convolutional Kernel Modulator (CKM)
Input: $x\in\mathbb{R}^{B\times H\times W\times T\times C}$
Output: $\hat{x}\in\mathbb{R}^{B\times H\times W\times 1\times C}$
wbase: base kernel weights, B: resizing matrix
Step 1: Sample patch size
  • Sample patch size $k\in\{4,8,16\}$ once per forward pass.
  • If the base encoder/decoder to be flexified is multi-staged, split $k$ across stages.
Step 2: Encoding
  1. 1.
    For each encoder stage with assigned $k$:
    • Resize kernel: $w\leftarrow B^{T\dagger}w^{\text{base}}$
    • $x\leftarrow\texttt{Conv}(x,w,\texttt{stride}=k)$
Step 3: Transformer processing
  • Pass tokens through any transformer processor (architecture-agnostic)
Step 4: Decoding
  1. 1.
    For each decoder stage with assigned $k$:
    • Resize kernel: $w\leftarrow B^{T\dagger}w^{\text{base}}$
    • $\hat{x}\leftarrow\texttt{ConvTranspose}(x,w,\texttt{stride}=k)$
return $\hat{x}$
A.6 CSM pseudocode
Algorithm 2 Convolutional Stride Modulator (CSM)
Input: $x\in\mathbb{R}^{B\times H\times W\times T\times C}$
Output: $\hat{x}\in\mathbb{R}^{B\times H\times W\times 1\times C}$
$k^{\text{base}}$: fixed kernel size
Step 1: Padding
  • Pad $x$ with learned tokens based on boundary conditions.
Step 2: Stride selection
  • Sample stride(s) $s$ from {4, 8, 16}.
  • If the encoder/decoder is multi-stage, split $s$ across stages.
Step 3: Encoding
  1. 1.
    For each encoder stage with assigned $s$:
    • $x\leftarrow\texttt{Conv}(x,k^{\text{base}},\texttt{stride}=s)$
Step 4: Transformer processing
  • Pass tokens through transformer processor (architecture-agnostic)
Step 5: Decoding
  1. 1.
    For each decoder stage with assigned $s$:
    • $\hat{x}\leftarrow\texttt{ConvTranspose}(x,k^{\text{base}},\texttt{stride}=s)$
return $\hat{x}$
Appendix B Experiment details
B.1 Datasets
The datasets that we benchmarked are taken from The Well collection [20]. We selected 2D and 3D datasets with complex dynamics, ranging from biology (active_matter [78]) to astrophysics (supernova_explosion &\&& turbulence_gravity_cooling [79, 80];  turbulent_radiative_layer_2D [81]) and fluid dynamics (rayleigh_benard &\&&  shear_flow [82]) . Here is information about their resolution and physical fields (equivalent to a number of channels).
Dataset Resolution # of fields/channels
active_matter $256\times 256$ 11
rayleigh_benard $512\times 128$ 4
shear_flow $128\times 256$ 4
supernova_explosion $64\times 64\times 64$ 6
turbulence_gravity_cooling $64\times 64\times 64$ 6
turbulent_radiative_layer_2D $128\times 384$ 5
Table 4: Specifics of the datasets benchmarked.
B.2 Model configuration
The core transformer processor consists of multiple stacked processing blocks, each composed of three key operations: (1) temporal self-attention, which captures dependencies across time steps, (2) spatial self-attention, which extracts spatial correlations within each frame, and (3) a multi-layer perceptron (MLP) for feature transformation. The number of such processing blocks varies across model scales, as detailed in Table 5. The full configuration is vizualized in Fig. 6.
Table 5: Details of the various model architectures and scales explored.
Model Embedding Dim. MLP Dim. # Heads # Blocks Base down/up sampling rate of patch encoder/decoder
7.6M 192 768 3 12 [16, 16]
25M 384 1536 6 12 [16, 16]
100M 768 3072 12 12 [16, 16]
Figure 6: This visualization represents the overall architecture of our models. The CKM/CSM/Fixed patch are various strategies for the patch encoding and decoding. The patch encoder block is followed by $N$ transformer based spatio-temporal blocks containing sequential time attention, followed by full spatial attention, followed by MLP. The embedding dimension column in table 4 represents the hidden dimension that the CKM/CSM/Fixed patch strategies embed the real space patches into.
B.3 Training configuration - Axial ViT and Vanilla ViT experiments
In the following we list the training parameters and other details:
  1. 1.
    Batch size is set at 2 and the model and we train on 8 GPUs using PyTorch FSDP on A100-80Gb GPUs. No gradient accumulation
  2. 2.
    Adam Optimizer. learning rate of $10^{-4}$, weight decay of $10^{-4}$. We have hyperparameter tuned our learning rates amongst $10^{-4},10^{-3},10^{-2}$.
  3. 3.
    All models and datasets were trained using Normalized Mean Squared Error loss averaged over fields and space during training.
  4. 4.
    Prediction type: delta. “delta” predicts the change in the field from the previous timestep. This only affects training since validation and test losses are computed on reconstructed fields.
  5. 5.
    Position embedding: RoPE [83]. Drop path: 0.1. Epoch size: 100.
  6. 6.
    The Well dataset[20] comes with a train/valid/test set split, and we use their split for our experiments also. Within each dataset, for each set of simulation parameters, an 80/10/10 split is applied along the initial conditions. For example, if we have 100 initial conditions for each of 5 simulation parameters and capture 200 steps per simulation, we include 80 of these trajectories of 200 steps per simulation parameter in the training set, 10 in validation, and 10 in test.
  7. 7.
    As mentioned in the checklist, we have provided code for the main architectural blocks utilized in this paper, i.e. the encoder and decoder blocks of CSM and CKM that were used for flexifying axial and vanilla ViT architectures.
B.4 VRMSE metric
The variance scaled mean squared error (VMSE): it is the MSE normalized by the variance of the truth. It has been used for the benchmarking of the Well [20].
$\mathrm{VMSE}(u,v)=\frac{\langle\lvert u-v\rvert^{2}\rangle}{(\langle\lvert u-\bar{u}\rvert^{2}\rangle+\epsilon)}.$
We chose to report its square root variant, the VRMSE:
$\mathrm{VRMSE}(u,v)=\frac{\langle\lvert u-v\rvert^{2}\rangle^{1/2}}{(\langle\lvert u-\bar{u}\rvert^{2}\rangle+\epsilon)^{1/2}}.$
Note that, since $\mathrm{VRMSE}(u,\bar{u})\approx 1$, having $\mathrm{VRMSE}>1$ indicates worse results than an accurate estimation of the spatial mean $\bar{u}$.
Appendix C Binned spectral error analysis
C.1 Definitions
We investigate another metric for the robustness of our results, namely the binned spectral mean squared error (BSMSE). BSNMSE, introduced in [20], quantifies the mean squared error after applying a bandpass filter to the input fields over a specific frequency band $\mathcal{B}$. It is defined as:
$\mathrm{BSMSE}_{\mathcal{B}}(u,v)=\langle\lvert u_{\mathcal{B}}-v_{\mathcal{B}}\rvert^{2}\rangle,$ (18)
where the bandpass-filtered field $u_{\mathcal{B}}$ is given by:
$u_{\mathcal{B}}=\mathcal{F}^{-1}\left[\mathcal{F}\left[u\right]\mathbf{1}_{\mathcal{B}}\right].$ (19)
Here, $\mathcal{F}$ represents the discrete Fourier transform, and $\mathbf{1}_{\mathcal{B}}$ is an indicator function that selects frequencies within the band $\mathcal{B}$.
For each dataset, we define three disjoint frequency bands, $\mathcal{B}_{1}$, $\mathcal{B}_{2}$, and $\mathcal{B}_{3}$, which correspond to low, intermediate, and high spatial frequencies, respectively. These bands are determined by partitioning the wavenumber magnitudes evenly on a logarithmic scale.
The binned spectral normalized mean squared error (BSNMSE) is a normalized variant of the BSMSE metric, adjusting for the energy within each frequency band:
$\mathrm{BSNMSE}_{\mathcal{B}}(u,v)=\frac{\langle\lvert u_{\mathcal{B}}-v_{\mathcal{B}}\rvert^{2}\rangle}{\langle|v_{\mathcal{B}}|^{2}\rangle}.$ (20)
A $\mathrm{BSNMSE}$ value of 1 or greater indicates that the model performs worse than simply predicting zero coefficients at that scale.
C.2 BSNMSE for different datasets
Fig. 7 shows the next-step prediction BSNMSE of the test set FOR rayleigh_benard, turbulent_radiative_layer_2D, shear_flow  and supernova_explosion. The spectrum is divided into three frequency bins: low, mid and high being the smallest, intermediate and largest frequency bins whose boundaries are evenly distributed in log space. Crucially, as also emphasized in the main text, we note that the fixed patch models (green) are separately trained models, while CKM (orange) and CSM (blue) (here shown is the flexified vanilla ViT model) are single models trained at randomized patch/strides.
Following our results in Tables 2 and 3 of the main text, our story is reflected in the BSNMSE metric as well. In other words, it is clear from the BSNMSE plot that CSM/CKM (coupled with vanilla ViT for this plot) enable the creation of a single flexible model without any loss in accuracy (significant accuracy gain in fact for a range of cases). This enables the creation of a new class of these compute aware models for the PDE surrogate literature.
Figure 7: BSNMSE of the next-step prediction of the test set in the rayleigh_benard, turbulent_radiative_layer_2D, shear_flow  and supernova_explosion datasets. The spectrum is divided into three frequency bins: low, mid and high being the smallest, intermediate and largest frequency bins whose boundaries are evenly distributed in log space. Crucially, as also emphasized in the main text, we note that the fixed patch models (green) are separately trained models, while CKM (orange) and CSM (blue) (here shown is the flexified vanilla ViT model) are single models trained at randomized patch/stride sizes.
Appendix D CViT experiments
CViT experiments. We further evaluate our flexible patching strategy within the Continuous Vision Transformer (CViT) architecture [63], a recent vision transformer-based model tailored for PDE surrogate modeling. CViT features a patch-based transformer encoder coupled with a novel, grid-based continuous decoder that directly queries spatiotemporal locations. As the decoder operates without patch-based tokenization, we apply our kernel modulation (CKM) only to the encoder, leaving the rest of the architecture unchanged.
A flexified CViT therefore refers to the encoder being a CKM. For CViT [63], whose original design features a single-stage encoder (instead of multi-stage hMLP we used for our vanilla ViT and axial ViT experiments) and a decoder based on grid-based querying, CKM is applied only to the encoder, for the single stage; because the idea is to see if flexification on top of the base fixed patch architecture helps or not. These use cases demonstrate CKM’s modularity and adaptability: it can flexify arbitrary convolutional patching pipelines, without requiring architectural changes or any overengineering to the base convolutional or attention mechanism or task head.
Since CViT is designed for 2D problems, we restrict our evaluation to 2D datasets, ensuring a fair and focused test of patch flexibility without altering any other model components. This experiment demonstrates that our flexible tokenization strategy is architecture-agnostic and can be modularly integrated into advanced PDE models like CViT.
Table 6 shows that a flexified CViT consistently outperforms fixed patch size CViT models across four diverse 2D PDE datasets. As in earlier experiments, reducing patch size (e.g., from 16 to 4) increases token count and compute, but yields significantly improved accuracy. Importantly, a flexified CViT enables this accuracy-compute trade-off at inference without retraining, demonstrating the modularity and architectural agnosticism of our approach. This result also underscores how adaptive patching can be extended to complex hybrid architectures like CViT, unlocking new modes of inference-time flexibility in scientific surrogate modeling.
Dataset # Tokens/p.s. Flexified CViT f.p.s. CViT
shear_flow 2048/4 0.120 0.184
128/16 0.127 0.319
turbulent_radiative_layer_2D 3072/4 0.302 0.299
192/16 0.340 0.364
active_matter 4096/4 0.054 0.110
256/16 0.065 0.127
Table 6: CViT next-step prediction performance across datasets and token counts, comparing CKM and Fixed patch size (f.p.s.). As before, the f.p.s. models are separately trained models while the flexible models are one single model capable of handling multiple patch resolutions at inference. Similar to the vanilla ViT and axial ViT architectures, we find that CViT consistently gains from smaller patch sizes, can be flexified for practical deployment and an improvement in accuracy.
In terms of rollouts, since CViT uses a query based continuous decoder, instead of a patch based decoder, it doesn’t inherently suffer from the issues of patch artifacts. Even for that case, Table 7 shows that flexified CViT performs a lot better than the corresponding static patch version. This again demonstrates that our flexification approach is useful even for cases of hybrid encoder-decoder architectures where only the encoder is patch based, but the decoder is not. This continues to highlight the architecture agnostic–ness and the SOTA performance of our flexible models.
Dataset Flexified CViT f.p.s. 16 CViT $\Delta$
shear_flow 0.063 0.2598 +75.7%
turbulent_radiative_layer_2D 0.540 0.568 +4.9%
active_matter 0.78 3.33 +76.57%
Table 7: 10-step rollout VRMSE comparison across datasets for CViT vs fixed patch size 16 (f.p.s) models. $\Delta$ indicates the best percentage improvement (i.e., lowest VRMSE vs. fixed patch baseline).
Model details for CViT: For the base CViT architecture, we utilized the CViT-S configuration in [63] (See Table 1 of this paper), setting: Encoder layers = 5, Embedding dim = 384, MLP width = 384, Heads = 6. The embedding grid size for CViT was set to be the same as the dataset size.
CViT Patch Embeddings.
The ViT encoder takes as input a gridded representation of the input function $u$, yielding a spatio-temporal data tensor $\mathbf{u}\in\mathrm{R}^{T\times H\times W\times D}$ with $D$ channels. The input is patchified into tokens $\mathbf{u}_{p}\in\mathrm{R}^{T\times\frac{H}{P}\times\frac{W}{P}\times C}$ by tokenizing each 2D spatial frame independently, following the process used in standard Vision Transformers [9]. The patching process involves a single convolutional downsampling layer. Trainable 1D temporal and 2D spatial positional embeddings are added to each token. These are absolute position encodings, unlike the RoPE encodings we used for vanilla and axial ViT experiments.
$\mathbf{u}_{pe}=\mathbf{u}_{p}+\text{PE}_{t}+\text{PE}_{s},\quad\text{PE}_{t}\in\mathrm{R}^{T\times 1\times 1\times C},\quad\text{PE}_{s}\in\mathrm{R}^{1\times\frac{H}{P}\times\frac{W}{P}\times C}.$ (21)
The above description of CViT has already been detailed in the original paper [63]. Our main experiments were to flexify this CViT architecture and show that under a given compute budget, training a flexible CViT is much more advantageous than training multiple fixed patch CViT models. To this end, we adapt CKM to flexify the convolutional layer of CViT, randomizing the patch size at the encoding stage. Additionally, since base CViT has absolute learned position embedding, we resized the position embedding at each step of the forward pass as well using bilinear interpolation to make it compatible with the randomized patch size.
Goals with flexified CViT: Our goal with the CViT experiments is to show that, similar to the axial ViT and vanilla ViT results, CViT also benefits from flexification. In other words, given a fixed compute budget, it is much more beneficial to train a flexible CViT model than to train and maintain multiple static resolution CViT models. This leads to the creation of a flexible CViT, similar to the flexible versions of axial and vanilla ViTs we created earlier.
First we show that similar to axial and vanilla ViT, CViT also consistently gains in accuracy as the patch size, aka token counts is decreased (Table 6). Additionally, we show that our alternating patch strategy for CViT leads to dramatically better accuracies than the fixed patch models (Table 7). Through these experiments, we enable the enable the flexification of this recent architecture. This experiment further demonstrates the architecture agnostic–ness of our methods; meaning flexification can be made compatible with a range of base architectures, as exemplified by our results in the main text, augmented the flexification of this advanced hybrid architecture like CViT. Flexified CViT is compute-adaptive meaning it can adapt to various downstream compute/accuracy requirements at inference after being trained only once, eliminating the need to train and maintain multiple models.
Training details for CViT experiments. We follow a consistent training protocol to ensure a fair comparison between fixed-patch and flexified CViT models. All models use the CViT-S architecture from [63], configured with 5 encoder layers, an embedding dimension of 384, MLP width 384, and 6 attention heads. The spatial embedding grid size is matched to the input resolution of each dataset.
We fix the learning rate at $10^{-4}$, chosen after a hyperparameter search over $\{10^{-2},10^{-3},10^{-4}\}$. Weight decay is set to $10^{-5}$. Batch size is chosen to fully utilize GPU memory for each dataset. Training is conducted on a single NVIDIA A100 80GB GPU.
Due to compute limitations, training epochs vary across datasets, but remain consistent within each dataset. Concretely:
  • For shear_flow, each fixed-patch model (patch sizes 4, 8, and 16) is trained for 30 epochs.
  • For active_matter, each is trained for 40 epochs.
  • For turbulent_radiative_layer_2D, each is trained for 160 epochs.
The corresponding flexified CViT models (i.e., with CKM) are trained using the combined compute budget of all three fixed-patch models. This allows a direct, compute-equivalent comparison. As described in Sec. 5.2 of the main paper, this setup addresses a practical question: under a fixed training budget, is it better to train multiple static models or one flexible model? If the flexible model enables training without any loss in validation accuracy, we have a flexible model by training only once eliminating the need to train and maintain multiple fixed patch models.
Figure 8: Rollout for the $v_{y}$ shear_flow  dataset for the CViT model. Top: Ground Truth (G.T.); Middle: Flexified CViT; Bottom: Fixed Patch 16 (F.P.S. 16) models.
Figure 9: Rollout for the density turbulent_radiative_layer_2D  dataset for the CViT model. Top: Ground Truth (G.T.); Middle: Flexified CViT; Bottom: Fixed Patch 16 (F.P.S. 16) models.
Appendix E Patch artifacts
Patch artifacts are widely present in patch based PDE surrogates. In this section, we will provide extended evidence for the existence of these artifacts. Note that these patch artifacts are a problem for the architectures where both the encoder and decoder are patch based, so this is not a problem for an architecture like CViT.
E.1 Our experiments
We have investigated and found these artifacts to present in a range of patch based ViT architectures. Below we will show patch artifacts arising in three different types of ViT based architectures, in different kinds of datasets: axial ViT [22], full ViT [9] and Swin [76].
Figs. 10 and 11 reflects the presence of these artifacts in rollouts of two example fields of the active_matter  and turbulence_gravity_cooling  datasets respectively. Fig. 12 also shows artifacts arising when the base attention is changed to swin. Note that having a shifted window attention does not change the fact that the encoder and decoder still patchify with fixed size patches in conventional ViTs. Therefore even though attention can happen in a shifted-window format this fundamental limination imposed by the encoding-decoding process still remains.
Figure 10: Artifacts arising in fixed patch models of the axial+ViT fixed patch 16 model. Results are shown for the $E_{xx}$ field of active_matter  dataset.
Figure 11: Artifacts arising in fixed patch models of the vanilla+ViT fixed patch 16 model. Results are shown for the $v_{x}$ field of turbulence_gravity_cooling  dataset.
Figure 12: Artifacts arising in a swin transformer model where the spatial module is replaced by a swin transformer, keeping everything else the same. Results are shown for the density field of turbulent_radiative_layer_2D  dataset.
We have additionally tested that these artifacts are a fundamental problem with the standard patch based ViTs, and they do not simply arise due to training details, such as different learning rates. Fig. 13 demonstrates that patch artifacts are agnostic to learning rates. This is expected because they arise because of the fundamental limitation of fixed patch grids.
Figure 13: Artifacts in the density field of turbulent_radiative_layer_2D  dataset at rollout step 20. Artifacts are agnostic to what learning rate is used for training.
We have additionally tested that these artifacts also arise in neural operator models like AFNO [74], which also has a patch based encoder and a decoder. We took the AFNO model provided as a baseline comparison in the Well benchmark suite [20], and trained it on the turbulent_radiative_layer_2Ddataset, and as expected due to its patch based encoding and decoding, patch artifacts arise here as well as shown in Fig. 14.
Figure 14: Artifacts in the density field of turbulent_radiative_layer_2D  dataset at rollout step 20. This is the AFNO [74] model. Patch artifacts also arise here because even though its a neural operator based method, but the encoder and decoder still uses ViT based patch encoding/decoding.
E.2 Other papers
These artifacts have also been seen in models where patch based encoders and decoders are used. For example, a recent study on pretraining on the dataset of the Well by Morel et al. [66]. The artifacts can be seen in Figures 10 (3rd and 4th row), Figure 11 (3rd and 7th row), Figure 12 (3rd and 7th row)and Figure 13 (3rd, 4th and 7th row). In this paper, these artifacts are most common in the Multiple Physics Pretraining, MPP [7] approach, which has both a patch based and a decoder.
The broad goal in this section is to demonstrate that these patch based artifacts are quite commonly seen in models that incorporate patch based encoding and decoding. Our flexible patching/striding methods can be incorporated into a wide variety of these models, and we anticipate that the rollout strategies we discussed in the main text can be improve the overall stability, accuracy and mitigate patch artifacts in a range of these models. Due to limited compute, experiments will every possible architecture is difficult, but our study opens up possibilities in exploring this direction.
Appendix F Baselines used and more rollouts
In section 4.1 of the main paper (Table 1), we provided baseline comparisons with respect to a range of SOTA models which are not patch based. On the neural operator side, we included FFNO [69]. We also test against a SOTA convolution based model, SineNet [70]. Additionally, we tested against a recent Physics Attention based scheme, Transolver [71]. In addition to these comparisons, we also perform comparisons against baseline models provided with the Well dataset module, and the well paper [20]. These models were TFNO [68], FNO [72], U-Net [84] and CNext-U-Net [73]. In Table 2, we reported the best baseline results out of these comparisons against the Well benchmarks. These non-patch baselines covered a large range of comparisons. Below we describe these models and the settings used to make comparisons:
SineNet:
This model consists of multiple sequentially connected U-shaped network blocks, referred to as waves [70]. While traditional U-Net architectures use skip connections to support multi-scale feature processing, the authors argue that forcing features to evolve across layers leads to temporal misalignment in these skip connections, ultimately limiting model performance. In contrast, SineNet progressively evolves high-resolution features across multiple stages, reducing feature misalignment within each stage and improving temporal coherence. This model has achieved SOTA performance on a range of complex PDE tasks. Refer to the original paper for more details.
We used their open source code from: https://github.com/divelab/AIRS/blob/main/OpenPDE/SineNet/pdearena/pdearena/modules/sinenet_dual.py. We use the SineNet-8-dual model with configurations outlined in their codebase, taken from https://github.com/divelab/AIRS/blob/main/OpenPDE/SineNet/pdearena/pdearena/models/registry.py (sinenet8-dual model). The network has 64 hidden channels with 8 waves. The learning rate is set to be $2\times 10^{-4}$, set after doing a coarse tuning amongst $10^{-4},2\times 10^{-4},5\times 10^{-4},10^{-3}$. The resulting model has around 35M parameters. Training was performed on a single Nvidia A100-80GB GPU. Batch size was set at 16.
F-FNO:
Fourier neural operators [72] perform convolutions in the frequency domain and were originally developed for PDE modeling. We compare to a state-of-the-art variant, the factorized FNO [69] which improves over the original FNO by learning features in the Fourier space in each dimension independently, a process called Fourier factorization. We used the open source FFNO codebase provided in the original publication. We used a small version of the FFNO model with 4 layers, 32 fourier modes and 96 channels. This amounted to a model size of $\sim$ 7M. During training, a learning rate of 10-3 is utilized after a parameter tuning in the range 10-4, 10-3 and 10-2. Training was performed on a single Nvidia A100-80GB GPU. Batch size was set to be 16.
Transolver:
In this model [71], the authors propose a new Physics-Attention scheme to adaptively split the discretized domain into a series of learnable slices of flexible shapes, where mesh points under similar physical states is ascribed to the same slice. The model was found to achieve SOTA performance in a range of benchmarks including large-scale industrial simulations, including car and airfoil designs. We use their open source codebase and adapt it to train on the Well. We used 8 hidden layers, 8 heads, 256 hidden channels, and 128 slices. This amounted to a model of size 11M. This configuration matches the configuration presented in their main experiments. Training was performed on a single Nvidia A100-80GB GPU. Batch size was set to be 16.
As shown in Table 1 of the main paper, transolver generally struggled with the PDE tasks we tested on. We tested by tuning the learning rates varying between 10-3, 10-4 and 10-2 without any significant improvement in performance. In the original paper, for autoregressive prediction tasks, they tested on $64\times 64$ Navier Stokes, and downsampled their Darcy data into $85\times 85$ resolution. Our datasets, on the other hand all exceed $128\times 128$, and therefore the training was much slower. In order to have a fair comparison, we do not downsample our datasets for training the transolver model. The transolver code we used, taken directly from their codebase can be seen in: https://github.com/xxzebracrossing/well_exps/tree/main/the_well/benchmark/models/transolver.
TFNO, FNO, U-Net, CNext-U-Net
: These additional baselines with default configs are provided together as a part of the Well benchmarking suite [20]. We use these default baselines. TFNO is provided by neuraloperator v0.3.0. Defaults are 16 modes, 4 blocks and 128 hidden size. U-net classic has default configs of: initial dimension = 48, spatial filter size = 3, block per stage = 1, up/down blocks = 4, bottleneck blocks = 1. CNext-U-Net has configs: spatial filter size = 7, initial dimension = 42, block per stage = 2, up/down blocks = 4 and bottleneck blocks = 1. Finally, the FNO model has 16 modes, 4 blocks and 128 hidden size. As noted before, these model configs are provided as a part of the Well baseline package and more details can be found there. In Table 2 of the main text, we showed the best benchmark results from the Well baseline suite and found that our models consistently outperformed these baselines. Table 8 shows the additional baselines.
Dataset FNO TFNO UNet (Classic) CNext-U-Net
active_matter 0.262 0.257 0.103 0.033
rayleigh_benard 0.355 0.300 0.469 0.224
shear_flow 0.104 0.110 0.259 0.105
supernova_explosion 0.378 0.379 0.306 0.318
turbulence_gravity_cooling 0.243 0.267 0.675 0.209
turbulent_radiative_layer_2D 0.500 0.501 0.241 0.226
Table 8: Next-step VRMSE across Well benchmark datasets using baselines provided in the Well.
All of the above mentioned baselines were trained with an epoch size = 100, batch size = 16, and number of epochs = 200. This is chosen so that the total number of observed samples is the same for our flexible models ensuring that the comparison is fair.
Out of all of these baselines, we found that F-FNO and SineNet gave the most competitive results to our models, as shown in Table 1 on two of our complex 2D datasets. Given the competitive performance of F-FNO and SineNet on the two datasets shown in the text, we performed additional comparisons of these baselines on more datasets shown in Table 9. Clearly, these additional comparisons tell the same story that out flexible architecture-agnostic models outperform SOTA architectures on complex physics tasks.
Model FFNO SineNet Vanilla ViT 7M (Ours) Axial ViT 45M (Ours)
Number of parameters 7M 35M CKM CSM CKM CSM
rayleigh_benard 0.94 $>1$ 0.192 0.137 0.215 0.1664
active_matter 0.567 0.760 0.522 0.563 0.390 0.384
Table 9: 10-step rollout VRMSE comparison across patch-free models and our patch-based surrogate variants. We report scores for both 7M and 100M parameter versions of CSM and CKM. Results are shown for vanilla and axial self-attention scheme.
Figure 15 visually shows that CKM/CSM + axial ViT models (45M size) are significantly better visually than the SineNet 35M model in long rollout predictions as well. These rollouts for a highly complex and chaotic rayleigh_benard  dataset. This once again highlights the state-of-the-art performance achieved by our flexible models.
Figure 15: Figure showing a rollout for the rayleigh_benard  dataset. From top to bottom: ground truth (G.T.), CKM + axial ViT, CSM + axial ViT, SineNet. They all have comparable sizes in the order of 40M parameters.
To further prove that the effectiveness of our flexible models, we show the 44th step rollout of all four fields of the highly complex and chaotic rayleigh_benarddataset in Figs. 16 for CSM + vanilla ViT, 17 for CKM + vanilla ViT and 18 for FFNO. We chose to show FFNO because it was found to be the best performing neural operator model in our baseline suite. All of the models have a comparable size of about 7M paramaters. Along with the metrics noted in Table 9, we find compelling visual evidence that CSM/CKM perform dramatically better than a competitive model like F-FNO at a similar parameter level. This further goes to show the strong performance of our approaches.
Figure 16: Figure showing the 44th step rollout of the 7M CSM + Vanilla ViT model.
Figure 17: Figure showing the 44th step rollout of the 7M CKM + Vanilla ViT model.
Figure 18: Figure showing the 44th step rollout of the 7M F-FNO model.
More rollouts.
Additionally, provide some more rollouts for different datasets for our flexible models. Figs. 19 and 20. Consistent with the message of our paper, we find that the CKM and CSM strategies help to stabilize and significantly improve rollouts compared to fixed patch models; along with beating a range of state-of-the art models that we have extensively discussed.
Figure 19: Rollouts of the active_matter  dataset.
Figure 20: Rollouts of the turbulence_gravity_cooling  dataset.
Appendix G Ablation studies
G.1 Model size ablation
For the shear_flow  dataset, we trained the CSM/CKM and fixed patch (for the vanilla ViT self-attention) models. The next step test set VRMSE are shown in Fig. 21. The results are to be interpreted in the same way as Fig. 3 in the main text; i.e. the green points are separately trained static patch models which are being evaluated at the corresponding training patch/size or token counts at inference. CSM and CKM on the other hand are a single model, which are being evaluated at different token counts at inference. The plots show that CKM/CSM outperform the corresponding static patch counterparts consistently at a range of parameter counts varying from 7M to 100M. This ablation shows that our story is consistent across scaling parameter count.
For further investigation, we repeated the ablation experiment for the smallest 7M model for our other benchmark datasets, shown in Table 10. We find CSM/CKM to be consistently performing competitively with their fixed patch counterparts across our benchmark datasets. This shows, again that flexification is valuable to be incorporated into current patch based PDE surrogates, since it enables a flexible choice of patching/striding parameter enabling tunability with various downstream tasks and enable compute adaptivity in PDE surrogates.
Figure 21: Comparison of next step test VRMSE as a function of the number of tokens at inference for convolutional kernel modulation models (CKM), convolutional stride modulation models (CSM), and individually trained fixed patch models (7.6M, 25M and 100M parameters) for the shear_flow dataset.
Dataset Patch size Test VRMSE 7M Vanilla ViT
CKM CSM Fixed patch size
turbulent_radiative_layer_2D 3072 0.187 0.180 0.190
768 0.210 0.210 0.210
192 0.254 0.245 0.26
active_matter 4096 0.0405 0.0400 0.0409
1024 0.046 0.050 0.043
256 0.070 0.066 0.050
rayleigh_benard 4096 0.046 0.044 0.061
1024 0.060 0.060 0.073
256 0.080 0.074 0.083
supernova_explosion 4096 0.258 0.270 0.261
512 0.328 0.343 0.337
64 0.364 0.370 0.367
turbulence_gravity_cooling 4096 0.096 0.102 0.100
512 0.138 0.152 0.133
64 0.169 0.169 0.164
Table 10: Test VRMSE for different patch sizes (or, token counts) at inference time for three different 2D datasets using 7.6M parameter models. These values are for next step prediction.
G.2 Underlying patch size ablation in CKM
We vary the underlying patch size of CKM (coupled with axial ViT). We use a sequence of base p.s. between 4, 8, 16 and 32. We tested on the turbulent_radiative_layer_2Ddataset.
As shown in Table 11, we find that all of the base patch sizes perform comparably to each other. Our conclusion on the adaptive compute, i.e., our claim that flexible models can enable flexibility in PDE surrogates without compromising on the accuracy hold at all base patch size choices.
Base p.s. Inference p.s=4 Inference p.s=8 Inference p.s=16
4 0.157 0.181 0.246
8 0.156 0.179 0.244
16 0.160 0.175 0.216
32 0.161 0.179 0.221
Table 11: Ablation study on the effect of the base patch size used in CKM, evaluated on the turbulent_radiative_layer_2D  dataset with an Axial ViT backbone. Values are for next-step prediction VRMSE.
G.3 Sensitivity to patch diversity during training
Next we performed an experiment to determine the effect of omitting patch sizes during training. To this end, we train on patch/stride sizes of 4 and 16 for our flexible models, and then during inference, we test on patch/stride of 4, 8 and 16. Clearly, as shown in Table 12, when the evaluation is done on the omitted patch size of 8, the performance degrades for both CKM and CSM. This behavior is expected, and underlines the importance of training on a diverse patch/stride sizes to obtain the best generalization performance at inference.
Inference p.s.=4 Inference p.s.=8 Inference p.s.=16
CKM + axial ViT 0.158 0.453 0.230
CSM + axial ViT 0.160 0.377 0.222
Table 12: Performance drop when patch size 8 is omitted during training but used at inference. Models were trained only on patch sizes 4 and 16. Results highlight the importance of patch diversity during training to ensure generalization at inference patch/stride choices.
G.4 Alternating vs. random patch/stride rollout schedules
Next we experiment on another kind of rollout strategy – where we randomize the patches during rolling out temporally. We perform the experiment for both CKM and CSM on the vanilla ViT processor. The result is shown for two example datasets turbulent_radiative_layer_2D  and active_matter. We find that alternating schedule generally performs better than randomized schedules.
Dataset Model Alternating Randomized
active_matter CSM + vanilla ViT 0.359 0.364
CKM + axial ViT 0.351 0.394
turbulent_radiative_layer_2D CSM + vanilla ViT 0.370 0.370
CKM + vanilla ViT 0.409 0.426
turbulence_gravity_cooling CSM + vanilla ViT 0.559 0.625
CKM + vanilla ViT 0.527 0.597
Table 13: Rollout performance across datasets comparing alternating vs. randomized patch/stride schedules.
Appendix H Neural operator based methods benefitting from smaller patch sizes
To show that patch size is a critical parameter for some neural operator based methods as well, we conducted a preliminary experiment of the AFNO [74] baseline provided with the Well package. We took the turbulent_radiative_layer_2D  dataset and trained the AFNO model with a patch size of 16 and 4. Overall, we found that reducing AFNO’s patch size from 16 to 4 improves test VRMSE from 0.298 to 0.211 on the this dataset. Additionally, patch size was found to be a critical parameter in a recent DPOT [75] model which performed large-scale pretraining on PDEs.
Overall, these show that patch size is an important parameter in setups other than the ViT frameworks we considered in this paper. These findings suggest our flexible models can be valuable for extending compute-adaptive capabilities to a broader class of neural operator architectures, enabling task-aware inference-time control in both pretrained and task-specific surrogate models.